# Orbital-Selective Band Structure Evolution in BaFe2-xMxAs2 (M = Cr, Co, Cu, Ru and Mn) Probed by Polarization-Dependent ARPES

### v1

@@ Abstract
We present a systematic study of the evolution of the band structure in the Fe-based superconductor family BaFe2-xMxAs2 (M = Co, Cu, Cr, Ru and Mn) using polarization-dependent angle-resolved photoemission spectroscopy (ARPES). Low-substituted samples, with comparable spin-density wave transition temperatures ($T_{\text{SDW}}$), were chosen to facilitate controlled comparisons. The sizes of the central hole pockets ($\alpha$, $\beta$, and $\gamma$) remain largely unchanged across different substitutions, showing no clear correlation with either $T_{\text{SDW}}$ or the As height relative to the Fe planes. However, subtle trends are observed: a modest increase in the size of the $\eta_{\text{X}}$ electron pocket correlates with the suppression of $T_{\text{SDW}}$. Furthermore, the contraction of the $\eta_{\text{X}}$ pocket appears to be linked to an increase in the As height relative to the Fe planes. Our results suggest that the suppression of $T_{\text{SDW}}$ is primarily driven by changes in the Fe–As bond length, with the effect being more pronounced in electronic states with planar character. These findings provide insight into the electronic structure of BaFe2-xMxAs2.
@@ I Introduction
Current hypotheses in the literature suggest that structural distortions within the FeAs tetrahedra play a critical role in modifying the physical properties of these materials, potentially driving the system into a superconducting state. Specifically, it has been proposed that an optimal distance between the As atoms and the Fe plane, along with a specific Fe–As–Fe bond angle, are crucial factors favoring the emergence of superconductivity [8, 9, 10]. Lee et al.[11] were the first to report that an Fe–As–Fe bond angle of approximately $109.5^{\circ}$ favors high $T_{\text{SC}}$ values. Later, Mizuguchi et al.[12] showed that an optimal As height relative to the Fe planes of approximately $1.38,\text{\AA }$ also promotes high $T_{\text{SC}}$.
[…]
Yin et al. [22] calculated the theoretical FS for various FeSC compounds. For the BFA parent compound, the FS obtained using Density Function Theory plus Dinamical Mean Field Theory (DFT+DMFT) closely matches the experimental FS observed in our measurements, suggesting that a more strongly correlated scenario is favored in the BaFe2-xMxAs2 family. Their calculations predict the existence of three hole pockets at the center of the Brillouin zone (BZ): the outer pocket shows predominant $d_{xy}$ orbital character, while the two inner pockets exhibit a mixed contribution from the $d_{xz}$ and $d_{yz}$ orbitals. For the electron pockets, they predict two crossing ellipsoidal structures, each with mixed $d_{xy}$, $d_{xz}$, and $d_{yz}$ orbital character. The outer regions of the ellipses are dominated by $d_{xy}$, whereas the inner regions are more influenced by $d_{xz}$ and $d_{yz}$ orbitals.
Zhang et al. [25] reported ARPES measurements on optimally Co-substituted BFA, where similar features were observed, although with notable differences. In their results, the outer hole pocket exhibits mixed $d_{xy}$ and $d_{x^{2}-y^{2}}$ orbital contributions, while the electron pockets deviate from the previously predicted ellipsoidal shapes. Instead, they found an outer superellipse-like electron pocket (mainly $d_{x^{2}-y^{2}}$) and an inner pocket (primarily $d_{xz}$ and $d_{yz}$). The outer electron pocket originates from a deeper electron band, while the inner pocket is formed by a shallower one. Additionally, in the optimally Co-substituted compound, both inner hole bands lie below the Fermi level and do not contribute to the FS, leaving only the outer hole band crossing $E_{F}$.
Similar findings were reported by Pfau et al. [20] in their ARPES study of twinned and detwinned BFA. They observed that nematicity induces a splitting between the $d_{xz}$ and $d_{yz}$ orbitals and leads to renormalization of the low-energy band structure.
@@
In addition, we examine the evolution of the scattering rate extracted from ARPES spectra across the entire BaFe2-xMxAs2 series. A particular focus is given to its energy dependence, as our previous studies on Mn- and Cr-substituted compounds [26, 27], as well as work by Fink et al. [28], suggest a linear behavior characteristic of a marginal Fermi liquid. Our goal was to determine whether this behavior persists across different elements and to what extent it reflects the underlying electronic correlations in these systems.
@@ III Results and Discussion
Figure 2 presents a momentum distribution curve (MDC) waterfall plot with their specific fittings for the BFA parent compound. Two different fitting approaches were employed, both based on MDCs fitting, which are horizontal cuts at constant energy in the BMs. In the first approach, each band is modeled as a pair of Lorentzian peaks in the MDC spectra, with mirroed positions and equal widths. This method, widely used in ARPES analysis [31, 33, 34, 35], approximates the spectral function as a sum of Lorentzian peaks plus a background term (typically constant or linear). Due to the large number of fitting parameters involved, our initial fits fixed the peak positions to those obtained from the second derivative/curvature analysis. Once other parameters were determined, we performed a second fitting step in which the peak positions were allowed to vary while keeping some of the previously fitted parameters fixed.
@@
In the second approach, we attempted to directly fit the ARPES spectral function to the data. The spectral function is defined as:
@@
Figure 3: Summary of the ARPES measurements for the BFA parent compound. The upper panels (a1 and a2) correspond to measurements performed with $\pi$ polarization, while the lower panels (b1 and b2) correspond to those taken with $\sigma$ polarization. Left panels (a1 and b1) show the normal BMs, and right panels (a2 and b2) display the corresponding curvature BMs. Dashed lines serve as guide-to-eye to indicate the band dispersions. The colors red, green, and blue denote the $\alpha$, $\beta$, and $\gamma$ bands, respectively. Furthermore, at the X corner, the two electron bands $\delta$ (pink) and $\eta$ (orange) are indicated with dashed lines. The open squares denotes the points extracted via fitting or second derivative/curvature analysis.
Figure 3 presents the normal BMs (right panels) alongside their corresponding curvature BMs (left panels) for the BFA compound in the four possible experimental configurations: along the $\Gamma\text{X}$ and $\Gamma\text{M}$ directions, with $\pi$ and $\sigma$ polarizations. Each band of interest is labeled using distinct colors and names: red for the $\alpha$ band, green for the $\beta$ band, blue for the $\gamma$ band, pink for the $\delta$ band and orange for the $\eta$ band.
@@
Figure 5: Summary of the hole band dispersion analysis for the BaFe2-xMxAs2 family. Upper panels (a1 to f1) show data taken along the $\Gamma\text{X}$ direction, while lower panels (a2 to f2) show data taken along the $\Gamma\text{M}$ direction, both using $\pi$ and $\sigma$ polarizations. Circles represent data obtained with $\pi$ polarization, and squares with $\sigma$ polarization. Open markers (circles or squares) correspond to points extracted from curvature or second-derivative BMs, while filled markers indicate values obtained from MDC fitting of the normal BMs. Dashed lines are guide-to-eye of the band dispersions for each compound.
@@
In the figure 5, filled circles and squares represent band positions obtained by fitting the MDCs of the normal data. Open circles and squares correspond to positions determined from the minima of the MDCs in the curvature or second derivative data. Data measured with $\pi$ polarization are shown as circles (right side of the graphics), while data from $\sigma$ polarization are shown as squares (left side of the graphics).
@@
Next, we examine the evolution of hole and electron pocket sizes in relation to certain macroscopic properties of the studied materials. Figures 9(a) and (b) illustrate the evolution of the central hole pockets ($\alpha$, $\beta$, and $\gamma$), measured along the $\Gamma\text{X}$ direction (represented by filled circles and squares in light colors) and the $\Gamma\text{M}$ direction (open circles and squares in dark colors). Figures 9(c) and (d), in turn, show the evolution of the electron pockets ($\delta$ and $\eta$), measured along the $\Gamma\text{X}$ direction (filled markers, light colors) and the ZY direction (open markers, dark colors). In the left-hand panels, the samples are sorted by their SDW magnetic transition temperature ($T_{\mathrm{SDW}}$), while in the right-hand panels, they are sorted by the As height relative to the Fe planes. The latter quantity was extracted from Rietveld refinements of the x-ray diffraction data.
@@
Turning to the electron pockets shown in figures 9(c) and (d), we observe in panel (c), where the samples are sorted by $T_{\mathrm{SDW}}$, a slight trend: the $\eta_{X}$ electron pocket (closed orange circles) tends to increase in size as $T_{\mathrm{SDW}}$ decreases (shaded area in green and orange). However, the most prominent correlation appears in panel (d), where the samples are sorted by As height. Here, a shaded region highlights a clear relationship: as the As height increases, the size of the $\eta_{X}$ electron pocket decreases. Since the $\eta_{X}$ band is primarily associated with the $d_{xy}$ orbital, this observation suggests a link between structural distortions, suppression of $T_{\mathrm{SDW}}$, and modifications in the planar orbitals in the FS.
@@ V Acknowledgments
We acknowledge the Canadian Light Source (CLS) for time on Quantum Materials Spectroscopy Centre (QMSC) beamline under Proposal No. 014139 and support of Sergey Gorovikov. K.R.P acknowledge funding from the CNPq (grant 162327/2017-0) and CAPES (Finance Code 001); K.R.P., A.P.M., M.M.P., G.S.F., H.B.P., P.G.P. and C.A. acknowledge financial support from FAPESP (grant 2017/10581-1).

### v2

@@ Abstract
We present a systematic study of the evolution of the band structure in the Fe-based superconductor family BaFe2-xMxAs2 (M = Co, Cu, Cr, Ru and Mn) using polarization-dependent angle-resolved photoemission spectroscopy (ARPES). Low-substituted samples, with comparable spin-density wave transition temperatures ($T_{\text{SDW}}$), were chosen to enable controlled comparisons. The sizes of the central hole pockets ($\alpha$, $\beta$, and $\gamma$) remain largely unchanged across different substitutions, showing no clear correlation with either $T_{\text{SDW}}$ or the As height relative to the Fe planes. However, subtle trends are observed: a modest increase of the electron pocket with planar character that correlates with $T_{\text{SDW}}$ suppression, while its contraction appears linked to an increase in the As height relative to the Fe planes. Our results suggest that the suppression of $T_{\text{SDW}}$ is primarily driven by changes in the Fe–As bond length, with the effect being more pronounced in electronic states with planar character. These findings provide insight into the electronic structure of BaFe2-xMxAs2.
@@ I Introduction
Current hypotheses in the literature suggest that structural distortions within the FeAs tetrahedra play a critical role in modifying the physical properties of these materials, potentially driving the system into a superconducting state. Specifically, it has been proposed that an optimal distance between the As atoms and the Fe plane, along with a specific Fe–As–Fe bond angle, are crucial factors favoring the emergence of superconductivity [8, 9, 10]. Lee et al.[11] were the first to report that an Fe–As–Fe bond angle of approximately $109.5^{\circ}$ favors high superconducting critical temperature ($T_{\text{SC}}$) values. Later, Mizuguchi et al.[12] showed that an optimal As height relative to the Fe planes of approximately $1.38,\text{\AA }$ also promotes high $T_{\text{SC}}$.
This sensitivity of the physical properties to structural distortions of the FeAs tetrahedra is intimately connected to changes in the orbital contributions to the bands forming the Fermi surface (FS) in these materials. In particular, spectroscopic and theoretical studies have shown that the density of states (DOS) of the FeSCs near the Fermi level ($E_{F}$) is almost entirely derived from electrons in the five Fe $3d$ orbitals [13, 14, 15, 16, 17, 18]. Due to the crystal-field environment of the FeAs tetrahedra, the degeneracy of these five $3d$ orbitals is lifted into three nearly degenerate $t_{2g}$ states ($d_{xy}$, $d_{xz}$, and $d_{yz}$) and two $e_{g}$ states ($d_{x^{2}-y^{2}}$ and $d_{z^{2}}$). Consequently, even small perturbations to the tetrahedral geometry can induce significant changes in the orbital makeup of the FS and, therefore, modify the physical properties of the FeSCs.
Subsequent studies have proposed that an enhanced population of planar-electron orbitals (primarily $d_{xy}$ and $d_{x^{2}-y^{2}}$) at the FS may play a crucial role in promoting high-temperature superconductivity (HTSC) [19, 5]. The sensitivity of HTSC to the details of the electronic structure reflects its close connection to charge and spin fluctuations, which remain central to our current understanding of the FeSCs. Various theoretical approaches emphasize either the itinerant or localized character of these fluctuations [20], leading to different interpretations of how chemical substitutions influence the electronic structure and related physical properties.
Several theoretical and experimental investigations have attempted to clarify how the Fe $3d$ electrons organize into the bands that constitute the FS of the FeSCs. Yin et al. [21], for example, demonstrated that in the BFA parent compound the FS comprises five pockets: three hole pockets centered at the $\Gamma$ point of the Brillouin zone (BZ) and two electron pockets at the X point. The FS obtained using density functional theory combined with dynamical mean-field theory (DFT+DMFT) closely matches the FS observed in our ARPES measurements, indicating that electronic correlations play an important role in the BaFe2-xMxAs2 family. These findings are consistent with other theoretical [22, 23, 24] and ARPES [25, 26, 27] studies, which report similar FS shape and orbital contribution in BFA and its substituted compounds.
Concerning the orbital contributions to the electron and hole pockets, theoretical works such as those by Yin et al. [21] and van Roekeghem et al. [22] generally suggest that the FS is dominated by $t_{2g}$ orbitals, whereas the $e_{g}$ orbitals lie well below $E_{F}$ or contribute only at specific regions of the BZ, such as near the Z point. Similar conclusions were reached by Pfau et al. [25] in their ARPES study of twinned and detwinned BFA, where nematicity was shown to induce a splitting between the $d_{xz}$ and $d_{yz}$ orbitals and to significantly renormalize the low-energy band structure.
On the other hand, several experimental results point to additional complexity. Zhang et al. [28], studying optimally Co-substituted BFA, reported that the outer hole pocket exhibits a mixed $d_{xy}$ and $d_{x^{2}-y^{2}}$ orbital character. They also found that both inner hole bands lie below $E_{F}$ and thus do not contribute to the FS, leaving only the outer hole band crossing the Fermi level. Similar orbital assignments were reported by Evtushinsky et al. [26] for K-substituted BFA, with the main difference being that in this case both inner hole pockets contribute to the FS. Our results are consistent with these previous studies regarding the orbital character of the hole and electron pockets, while also revealing substitution-dependent energy shifts of the inner hole bands.
@@
In addition, we examine the evolution of the scattering rate as a function of the binding energy extracted from ARPES spectra across the entire BaFe2-xMxAs2 series. A particular focus is given to its energy dependence, as our previous studies on Mn- and Cr-substituted compounds [29, 30], as well as work by Fink et al. [31], suggest a linear behavior characteristic of a marginal Fermi liquid. Our goal was to determine whether this behavior persists across different elements and to what extent it reflects the underlying electronic correlations in these systems.
@@ III Results and Discussion
Figure 2(a) and (b) presents momentum distribution curves (MDCs) in a waterfall plot with their specific fittings for the BFA parent compound. Two different fitting approaches were employed, both based on MDCs fitting, which are horizontal cuts at constant energy in the BMs. In the first approach (Figure 2(c)), each band is modeled as a pair of Lorentzian peaks in the MDC spectra, with mirroed positions and equal widths. This method, widely used in ARPES analysis [34, 36, 37, 38], approximates the spectral function as a sum of Lorentzian peaks plus a background term (typically constant or linear). Due to the large number of fitting parameters involved, our initial fits fixed the peak positions to those obtained from the second derivative/curvature analysis. Once other parameters were determined, we performed a second fitting step in which the peak positions were allowed to vary while keeping some of the previously fitted parameters fixed.
@@
In the second approach (Figure 2(d)), we attempted to directly fit the ARPES spectral function to the data. The spectral function is defined as:
@@
Figure 3: Summary of the ARPES measurements for the BFA parent compound. The upper panels (a1 and a2) correspond to measurements performed with $\pi$ polarization, while the lower panels (b1 and b2) correspond to those taken with $\sigma$ polarization. Left panels (a1 and b1) show the raw BMs, and right panels (a2 and b2) display the corresponding curvature BMs. Dashed lines serve as guide-to-eye to indicate the band dispersions. The colors red, green, and blue denote the $\alpha$, $\beta$, and $\gamma$ bands, respectively. Furthermore, at the X corner, the two electron bands $\delta$ (pink) and $\eta$ (orange) are indicated with dashed lines. The open squares denotes the points extracted via fitting or second derivative/curvature analysis.
Figure 3 presents the raw BMs (left panels) alongside their corresponding curvature BMs (right panels) for the BFA compound in the four possible experimental configurations: along the $\Gamma\text{X}$ and $\Gamma\text{M}$ directions, with $\pi$ and $\sigma$ polarizations. Each band of interest is labeled using distinct colors and names: red for the $\alpha$ band, green for the $\beta$ band, blue for the $\gamma$ band, pink for the $\delta$ band and orange for the $\eta$ band.
@@
Figure 5: Summary of the hole band dispersion analysis for the BaFe2-xMxAs2 family. Upper panels (a1 to f1) show data taken along the $\Gamma\text{X}$ direction, while lower panels (a2 to f2) show data taken along the $\Gamma\text{M}$ direction, both using $\pi$ and $\sigma$ polarizations. Circles represent data obtained with $\pi$ polarization, and squares with $\sigma$ polarization. Open markers (circles or squares) correspond to points extracted from curvature or second-derivative BMs, while filled markers indicate values obtained from MDC fitting of the raw BMs. Dashed lines are guide-to-eye of the band dispersions for each compound.
@@
In the figure 5, filled circles and squares represent band positions obtained by fitting the MDCs of the raw data. Open circles and squares correspond to positions determined from the minima of the MDCs in the curvature or second derivative data. Data measured with $\pi$ polarization are shown as circles (right side of the graphics), while data from $\sigma$ polarization are shown as squares (left side of the graphics).
@@
Next, we examine the evolution of hole and electron pocket sizes in relation to certain macroscopic properties of the studied materials. Figures 9(a) and (b) illustrate the evolution of the central hole pockets ($\alpha$, $\beta$, and $\gamma$), measured along the $\Gamma\text{X}$ direction (represented by filled circles and squares in light colors) and the $\Gamma\text{M}$ direction (open circles and squares in dark colors). Figures 9(c) and (d), in turn, show the evolution of the electron pockets ($\delta$ and $\eta$), measured along the $\Gamma\text{X}$ direction (filled markers, light colors) and the ZY direction (open markers, dark colors). In the left-hand panels, the samples are sorted by their SDW magnetic transition temperature ($T_{\mathrm{SDW}}$), while in the right-hand panels, they are sorted by the As height relative to the Fe planes. The latter quantity was derived from Rietveld refinements of the X-ray diffraction data, enabling a direct determination of the lattice parameters, Fe–As bond length, and Fe–As–Fe bond angle.
@@
Turning to the electron pockets shown in figures 9(c) and (d), we observe in panel (c), where the samples are sorted by $T_{\mathrm{SDW}}$, a slight trend: the $\eta_{X}$ electron pocket (closed orange circles) tends to increase in size as $T_{\mathrm{SDW}}$ decreases (shaded area in green and orange). However, the most prominent correlation appears in panel (d), where the samples are sorted by As height. Here, a shaded region appears to show a clearer relationship: as the As height increases, the size of the $\eta_{X}$ electron pocket decreases. Since the $\eta_{X}$ band is primarily associated with the $d_{xy}$ orbital, this observation suggests a link between structural distortions, suppression of $T_{\mathrm{SDW}}$, and modifications in the planar orbitals in the FS.
@@ V Acknowledgments
We acknowledge the Canadian Light Source (CLS) for time on Quantum Materials Spectroscopy Centre (QMSC) beamline under Proposal No. 014139 and support of Sergey Gorovikov. K.R.P acknowledge funding from the CNPq (grant 162327/2017-0) and CAPES (Finance Code 001); K.R.P., A.P.M., M.M.P., G.S.F., H.B.P., F.A.G., P.G.P. and C.A. acknowledge financial support from FAPESP (grant 2017/10581-1, 2023/17024-1, 2021/02763-8, 2018/08040-5 and 2024/13291-8) and CNPq (Grants No. 405408/2023-4 and 311783/2021-0). We acknowledge the support of the INCT project Advanced Quantum Materials, involving the Brazilian agencies CNPq (Grant 408766/2024-7), FAPESP, and CAPES.